A water tank has a hole 3.0 cm^2 in area located 2.0 m below the water surface. How many kilograms of water emerge from the hole per second?


1
Expert's answer
2021-08-12T06:20:03-0400
"\frac{dm}{dt}=\rho Av\\\frac{dm}{dt}=(3\cdot10^{-4})(1000)\sqrt{2(9.8)(2)}=1.9\frac{kg}{s}"